Langley Corporation currently collects all of its customer payments in Greensboro. By going to a new lockbox system with boxes in Chicago and Boulder, Langley expects to reduce the time from when customers mail their checks to when the funds are collected. It expects the time to fall from an average of 9 days to an average of 4 days, a 5-day savings. Langley collects $50,000 per day. The extra costs associated with the lockbox are $15,000 per year. Langley's opportunity cost of funds is 10% per year. What is the expected annual before-tax profit of using the new lockbox system?

Group of answer choices

$15,000

$10,000

$25,000

none of these
Answer #1

Savings of the loackbox system = Days saved * amount saved per day * Cost

= 5*50000*10%

= 25000

Annual Profit = Savings-cost

= 25000-15000 = $10000